In the Michelson-Morley experiment, the length of the paths of the two beams is 11 m each.



The wavelength of the light used is 6000 ̊A. If the expected fringe shift is 0.4, calculate the



velocity of the earth relative to the ether.

1
Expert's answer
2022-04-26T14:56:11-0400

"n=\frac{2lv^2}{\lambda c^2},\implies"

"v=c\sqrt{\frac{n\lambda}{2l}}=31.3~\frac{km}s ."